# A Fully Data-Driven Value Iteration for Stochastic LQR: Convergence, Robustness and Stability

### v2

@@ 1. Introduction
Notation: $\mathbb{R}$ (resp. $\mathbb{R}_{+}$) denotes the set of (nonnegative) real numbers; $\mathbb{Z}_{+}$ is the set of nonnegative integers; $\mathbb{S}^{n}$ is the set of real symmetric matrices of dimension $n$; $\mathbb{S}^{n}_{+}$ (resp. $\mathbb{S}^{n}_{++}$) is the set of real symmetric and positive semi-definite (definite) matrices of dimension $n$; $\otimes$ denotes the Kronecker product; $I_{n}$ denotes the identity matrix of dimension $n$ while $0_{m\times n}$ denotes the $m\times n$ zero matrix; $\|\cdot\|_{2}$ denotes the Euclidean norm for vectors and the spectral norm for matrices; for a function $\mathbb{F}\rightarrow\mathbb{R}^{n\times m}, u\|u\|_{\infty}$ denotes its $l^{\infty}$-norm when $\mathbb{F}=\mathbb{Z}_{+}$, and $L^{\infty}$-norm when $\mathbb{F}=\mathbb{R}$. $\mathbf{1}_{[0,T]}$ denotes an indicator function, that is $\mathbf{1}_{[0,T]}(t)=1$ if $t\in[0,T]$, and $\mathbf{1}_{[0,T]}(t)=0$ otherwise. For matrices $X\in\mathbb{R}^{m\times n},Y\in\mathbb{S}^{m}$, define $\mathrm{vec}(X)\triangleq\left[X_{1}^{\top},X_{2}^{\top},\ldots,X_{n}^{\top}\right]^{\top}\in\mathbb{R}^{nm}$, $\mathrm{svec}(Y)\triangleq\left[y_{11},\sqrt{2}y_{12},\ldots,\sqrt{2}y_{1m},y_{22},\sqrt{2}y_{23},\ldots,\sqrt{2}y_{m-1,m},y_{m,m}\right]^{\top}$ $\in\mathbb{R}^{\frac{1}{2}m(m+1)}$, where $X_{i}$ is the $i^{th}$ column of $X$. $\mathrm{vec}^{-1}(\cdot)$ and $\mathrm{svec}^{-1}(\cdot)$ are the operations such that $\mathrm{vec}^{-1}(\mathrm{vec}(X))=X$, and $\mathrm{svec}^{-1}(\mathrm{svec}(Y))=Y$. For $n\in\mathbb{Z}_{+}$, define $\tilde{n}\triangleq\frac{n(n+1)}{2}$. For a vector $v\in\mathbb{R}^{n}$, define $\tilde{v}\triangleq\mathrm{svec}(vv^{\top})\in\mathbb{R}^{\tilde{n}}$. For a vector $v\in\mathbb{R}^{n}$, define $\mathcal{E}(v)\triangleq\left[v_{1},\dots,v_{n-1}\right]^{\top}\in\mathbb{R}^{n-1}$ the truncation of the last element of $v$. For $Y_{1},Y_{2}\in\mathbb{S}^{m}$, $Y_{1}\preceq Y_{2}$ means that $Y_{2}-Y_{1}\in\mathbb{S}^{n}_{+}$. For $S\in\mathbb{S}^{n}_{+}$, $S^{\frac{1}{2}}$ denotes the upper triangular Cholesky factor of $S$. For a matrix $\in\mathbb{R}^{n\times n}, ZZ^{\dagger}$ denotes its Moore-Penrose pseudoinverse. For $Z\in\mathbb{R}^{m\times n}$, define $\mathcal{B}_{r}(Z)\triangleq\left\{X\in\mathbb{R}^{m\times n}\lVert X-Z\rVert_{1}<r\right\}$ to be the open ball of radius $r$ centered at $Z$, and $\overline{\mathcal{B}}_{r}(Z)$ its closure, similarly $\mathcal{B}^{\infty}_{r}(Z)$ define the ball for the infinity-norm.
@@ 2.2. The Stochastic Linear Quadratic Regulator Problem in Discrete-Time
In addition, $A-BK^{*}$ is Hurwitz, i.e. the spectral radius $\rho(A-BK^{*})<1$ (see, for example, [44, Section 2.4]). For any given $P\in\mathbb{S}_{+}^{n}$, the resulting control is
@@
$\mathcal{Q}\left(P\right)$ $\triangleq\left(\begin{array}[]{cc}A^{\top}PA+S&A^{\top}PB\\ B^{\top}PA&B^{\top}PB+R\end{array}\right)$ (11)
$=\left(\begin{array}[]{cc}\left[\mathcal{Q}(P)\right]_{xx}&\left[\mathcal{Q}(P)\right]_{ux}^{\top}\\ \left[\mathcal{Q}(P)\right]_{ux}&\left[\mathcal{Q}(P)\right]_{uu}\\ \end{array}\right)\in\mathbb{R}^{l\times l}$ (14)
@@ 2.3. Problem Formulation
In the setting of learning-based control, system dynamics are unknown and must be approximated using data-driven approaches. In some situations, the cost function may also be partially or completely unknown. Nonetheless, data employed in these approaches can be affected by disturbances such as system noise and numerical errors. Consequently, each iteration of the VI yields an approximate value matrix $P_{i+1}$. To incorporate errors and noise, we formulate the following inexact version of the VI.
@@ 3. Robust Least-Squares VI
$\mathbb{E}\left[X_{t+1}^{\top}PX_{t+1}+c_{t}|\mathcal{F}_{t}\right]$ $=(Ax_{t}+Bu_{t})^{\top}P(Ax_{t}+Bu_{t})+x_{t}^{\top}Sx_{t}+u_{t}^{\top}Ru_{t}$
$\qquad+\mathbb{E}\left[(C\epsilon_{t+1})^{\top}P(C\epsilon_{t+1})|\mathcal{F}_{t}\right]$ (29)
$=\left(\begin{array}[]{c}x_{t}\\ u_{t}\end{array}\right)^{\top}\mathcal{Q}(P)\left(\begin{array}[]{c}x_{t}\\ u_{t}\end{array}\right)+\mu(P)$ (34)
@@
For any $c>0$, there exists $T_{c}>0$, such that for all $T>T_{c}$
$\Theta>cI_{\tilde{l}+1}\textit{ with probability one.}$ (38)
@@
We add exploration noise $\eta_{t}\sim\mathcal{N}(0,\Sigma_{\eta}^{2}I)$ in the behavior policy $u_{t}$ so that collected data are linearly independent, ensuring that $\theta_{T}$ has full column rank and is invertible for $T$ large enough. This is consistent with the persistent excitation condition widely used in adaptive control [4, 60] and is similarly prevalent in various RL methods (see, for example, [28], [34], [36]).
@@
We now present the R-LSVI algorithm in Algorithm 1. The method operates entirely without access to system matrices $(A,B,S,R)$. We sample data by applying the behavior policy $u_{t}$ defined in (28) and observing $(x_{t},c_{t})$ along trajectories of (26),(27) and (5), from which we construct the data matrices (40). The subsequent steps follow Procedure 1, with $\hat{Q}_{i}$, computed via (39), replacing $Q_{i}$. Finally, we obtain the control gain $\hat{K}$ using formula (16), as it is the primary quantity of practical interest. Note that contrary to the PI Procedure, we do not need the knowledge of a stabilizing initial controller $K_{0}$.
@@
 Collect data $(x_{t},u_{t})$ from (26),(27) and (28).
@@ 4.1. Convergence of the Exact VI
The exact value iteration in Procedure 1 is exponentially asymptotically stable. That is, there exists a $\mathcal{KL}$-function $\beta$, such that $\forall i\in\mathbb{Z}_{+}\text{ and }\forall P_{0}\in\mathbb{S}_{+}^{n}$
@@ 4.3. Convergence of R-LSVI
$\lVert\hat{P}_{i}-P^{*}\rVert_{2}$ $\leq\beta(\lVert\hat{P}_{0}-P^{*}\rVert{2},i)+\gamma(\lVert\Delta\rVert_{\infty})$
@@
$\hat{Q}_{i}=\mathrm{svec}^{-1}(\mathcal{E}(\Theta^{\dagger}_{T}\Psi_{T}\mathrm{svec}(\hat{P}_{i})+\Theta^{-1}_{T}\Xi_{T}))\,.$ (64)
@@
Since $\hat{P}_{0}\in\mathcal{P}(\hat{P}_{0})$, by Proposition 5, $\left\|\Delta Q_{1}\right\|_{F}<d^{*}$ almost surely. By definition, $\hat{P}_{1}\in\mathcal{P}(\hat{P}_{0})$. Thus, by induction, we have $\left\|\Delta Q\right\|_{\infty}<d^{*}$ almost surely. By Theorem 4, the result follows. ∎
@@ 5. Empirical Results
  • Nominal Control [50, 18]: a plugin PI (VI) model-based procedure that first estimates the system and cost matrices $(A,B,Q,R)$ and then solves the stochastic LQR problem using Hewer’s PI [23] (Bertsekas’s VI see 1).
@@
This distinction matters for value iteration because the controller is derived from $\hat{P}$ (computed via the Lyapunov equation (74)) rather than from $\hat{P}{I{\text{max}}}$.
@@
We initialize each procedure with $x_{\text{init}}=0_{3}$ and $\hat{P}_{0}=\beta I_{3}$, with $\beta\sim\mathcal{U}\left(0,1\right)$ to ensure that $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$. Since the nominal PI, LSPI and O-LSPI are PI-based algorithms, it is necessary to select an initial control gain, $\hat{K}_{0}$. For consistency and fairness across comparisons, we choose $\hat{K}_{0}$ as the corresponding induced gains $\hat{K}_{0}=(R+B^{\top}\hat{P}_{0}B)^{-1}B^{\top}\hat{P}_{0}A$, as in [52]. For the policy gradient specifically, we take $\hat{K}_{0}=\alpha I_{3}$, with $\alpha\sim\mathcal{U}\left(0.1,1\right)$ stabilizing to enhance its convergence.
@@ 5.2. Adaptivity under Non-Quadratic Costs
$\tilde{c}(x_{t},u_{t})\triangleq x_{t}^{\top}Sx_{t}+(|u_{t}|^{\frac{\alpha}{2}})^{\top}R|u_{t}|^{\frac{\alpha}{2}}$ (76)
where $|\cdot|^{\frac{\alpha}{2}}$ is applied elementwise and $\alpha\in[1,3]$. Since this cost function is non-quadratic in $u$ for values other than $\alpha=2$, the optimal cost is not expected to adhere to the standard form $\mathrm{Tr}\left(C^{\top}P^{*}C\right)$. Consequently, direct comparisons of relative costs are not meaningful in this setting. Instead, we evaluate the cost obtained using a control input $u_{t}=-\hat{K}x_{t}$, where $\hat{K}$ is computed for each method using a sample size of $T=10^{5}$. The choice of $T$ is guided by prior results (Figure 1), which show that this sample size enables most methods to reach $100\%$ stability. To assess adaptability to cost misspecification independently of the non-stabilizing data-collection policy, we set $K_{c}=\alpha I_{3}$ with $\alpha\sim\mathcal{U}(0.1,0.2)$. In Figure 2, we observe that the ADP algorithms consistently outperform nominal model-based control in both adaptability and stability. In the left panel, for $\alpha=2$, the nominal control slightly outperforms the other methods, which is consistent with the exact model specification and the findings in Figure 1. However, beyond this point, R-LSVI systematically outperforms all other algorithms, exhibiting superior adaptability. Notably, the right panel shows that for $\alpha>2$, nominal model-based control becomes unstable, with stability dropping sharply below $20\%$. In contrast, non-model-based methods maintain robust performance, with R-LSVI achieving the highest stability, remaining above $80\%$ across all values of $\alpha$.
@@
Figure 2. Convergence and stability of the nominal control (VI and PI), LSPI, O-LSPI, R-LSVI and policy gradient algorithms on the data center cooling benchmark problem with non-quadratic cost. (Left) Costs as the exponent $\alpha$ varies. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.

### v3

@@ Abstract.
©2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 1. Introduction
Notation: $\mathbb{R}$ (resp. $\mathbb{R}_{+}$) denotes the set of (nonnegative) real numbers; $\mathbb{Z}_{+}$ is the set of nonnegative integers; $\mathbb{S}^{n}$ is the set of real symmetric matrices of dimension $n$; $\mathbb{S}^{n}_{+}$ (resp. $\mathbb{S}^{n}_{++}$) is the set of real symmetric and positive semi-definite (definite) matrices of dimension $n$; $\otimes$ denotes the Kronecker product; $I_{n}$ denotes the identity matrix of dimension $n$ while $0_{m\times n}$ denotes the $m\times n$ zero matrix; $\|\cdot\|_{2}$ denotes the Euclidean norm for vectors and the spectral norm for matrices; for a function $\mathbb{F}\rightarrow\mathbb{R}^{n\times m}, u\|u\|_{\infty}$ denotes its $l^{\infty}$-norm when $\mathbb{F}=\mathbb{Z}_{+}$, and $L^{\infty}$-norm when $\mathbb{F}=\mathbb{R}$. $\mathbf{1}_{[0,T]}$ denotes an indicator function, that is $\mathbf{1}_{[0,T]}(t)=1$ if $t\in[0,T]$, and $\mathbf{1}_{[0,T]}(t)=0$ otherwise. For matrices $X\in\mathbb{R}^{m\times n},Y\in\mathbb{S}^{m}$, define $\mathrm{vec}(X)\triangleq\left[X_{1}^{\top},X_{2}^{\top},\ldots,X_{n}^{\top}\right]^{\top}\in\mathbb{R}^{nm}$, $\mathrm{svec}(Y)\triangleq\left[y_{11},\sqrt{2}y_{12},\ldots,\sqrt{2}y_{1m},y_{22},\sqrt{2}y_{23},\ldots,\sqrt{2}y_{m-1,m},y_{m,m}\right]^{\top}$ $\in\mathbb{R}^{\frac{1}{2}m(m+1)}$, where $X_{i}$ is the $i^{th}$ column of $X$. $\mathrm{vec}^{-1}(\cdot)$ and $\mathrm{svec}^{-1}(\cdot)$ are the operations such that $\mathrm{vec}^{-1}(\mathrm{vec}(X))=X$, and $\mathrm{svec}^{-1}(\mathrm{svec}(Y))=Y$. For $n\in\mathbb{Z}_{+}$, define $\tilde{n}\triangleq\frac{n(n+1)}{2}$. For a vector $v\in\mathbb{R}^{n}$, define $\tilde{v}\triangleq\mathrm{svec}(vv^{\top})\in\mathbb{R}^{\tilde{n}}$. For a vector $v\in\mathbb{R}^{n}$, define $\mathcal{E}(v)\triangleq\left[v_{1},\dots,v_{n-1}\right]^{\top}\in\mathbb{R}^{n-1}$ the truncation of the last element of $v$. For $Y\in\mathbb{S}^{n}$, $\lambda_{\mathrm{min}}\left(Y\right)$ ($\lambda_{\mathrm{max}}\left(Y\right)$) denotes its smallest (largest) algebraic eigenvalue. For $Y_{1},Y_{2}\in{\mathbb{S}^{n}}$, $Y_{1}\preceq Y_{2}$ means that $Y_{2}-Y_{1}\in\mathbb{S}^{n}_{+}$. For $S\in\mathbb{S}^{n}_{+}$, $S^{\frac{1}{2}}$ denotes the upper triangular Cholesky factor of $S$. For a matrix $\in\mathbb{R}^{n\times n}, ZZ^{\dagger}$ denotes its Moore-Penrose pseudoinverse. For $Z\in\mathbb{R}^{m\times n}$, define $\mathcal{B}_{r}(Z)\triangleq\left\{X\in\mathbb{R}^{m\times n},\lVert X-Z\rVert_{{2}}<r\right\}$ to be the open ball of radius $r$ centered at $Z$, and $\overline{\mathcal{B}}_{r}(Z)$ its closure, similarly $\mathcal{B}^{\infty}_{r}(Z)$ defines the ball for the infinity-norm.
@@ 2.2. The Stochastic Linear Quadratic Regulator Problem in Discrete-Time
In addition, $A-BK^{*}$ is Schur stable, i.e. the spectral radius $\rho(A-BK^{*})<1$ (see, for example, [44, Section 2.4]). For any given $P\in\mathbb{S}_{+}^{n}$, the resulting control is
@@
$\mathcal{Q}\left(P\right)$ $\triangleq\left(\begin{array}[]{cc}A^{\top}PA+S&A^{\top}PB\\ B^{\top}PA&B^{\top}PB+R\end{array}\right)=\left(\begin{array}[]{cc}\left[\mathcal{Q}(P)\right]_{xx}&\left[\mathcal{Q}(P)\right]_{ux}^{\top}\\ \left[\mathcal{Q}(P)\right]_{ux}&\left[\mathcal{Q}(P)\right]_{uu}\\ \end{array}\right)$ (13)
@@ 2.3. Problem Formulation
In the setting of learning-based control, system dynamics are unknown and must be approximated using data-driven approaches. In some situations, the cost function may also be partially or completely unknown. Moreover, data employed in these approaches can be affected by disturbances such as system noise and numerical errors. Consequently, each iteration of the VI yields an approximate value matrix $\hat{P}_{i+1}$. To incorporate errors and noise, we formulate the following inexact version of the VI.
@@ 3. Robust Least-Squares VI
$\mathbb{E}\left[X_{t+1}^{\top}PX_{t+1}+c_{t}|\mathcal{F}_{t}\right]$ (28)
$\quad=(Ax_{t}+Bu_{t})^{\top}P(Ax_{t}+Bu_{t})+x_{t}^{\top}Sx_{t}+u_{t}^{\top}Ru_{t}$ (29)
$\quad\qquad+\mathbb{E}\left[(C\epsilon_{t+1})^{\top}P(C\epsilon_{t+1})|\mathcal{F}_{t}\right]$ (30)
$\quad=\left(\begin{array}[]{c}x_{t}\\ u_{t}\end{array}\right)^{\top}\mathcal{Q}(P)\left(\begin{array}[]{c}x_{t}\\ u_{t}\end{array}\right)+\mu(P)$ (35)
@@
There exists $c>0$, such that for $\phi_{t}\triangleq\frac{z_{t}}{\alpha_{t}}$ we have
$\Theta=\mathbb{E}_{\pi}[\phi_{t}\phi_{t}^{\top}]\succeq cI_{\tilde{l}+1}\,.$ (39)
@@
Assumption 2 is consistent with the persistent excitation condition widely used in adaptive control [4, 60] and is similarly prevalent in various RL methods (see, for example, [28], [34], [36]).
@@
We now present the R-LSVI algorithm in Algorithm 1. The method operates entirely without access to system matrices $(A,B,S,R)$. We sample data by applying the behavior policy $u_{t}$ defined in (27) and observing $(X_{t},x_{t},c_{t})$ along trajectories of (25),(26) and (5), from which we construct the data matrices (41). The subsequent steps follow Procedure 1, with $\hat{Q}_{i}$, computed via (40), replacing $Q_{i}$. Finally, we obtain the control gain $\hat{K}$ using formula (15), as it is the primary quantity of practical interest. Note that contrary to the PI Procedure, we do not need the knowledge of a stabilizing initial controller $K_{0}$.
@@
 Collect data $(X_{t},x_{t},u_{t})$ from (25),(26) and (27).
@@ 4.1. Convergence of the Exact VI
The exact value iteration in Procedure 1 is globally exponentially stable. That is, there exists a $\mathcal{KL}$-function $\beta$, such that $\forall i\in\mathbb{Z}_{+}\text{ and }\forall P_{0}\in\mathbb{S}_{+}^{n}$
@@ 4.3. Convergence of R-LSVI
$\lVert\hat{P}_{i}-P^{*}\rVert_{2}$ $\leq\beta(\lVert\hat{P}_{0}-P^{*}\rVert_{{2}},i)+\gamma(\lVert\Delta\rVert_{\infty})$
@@
$\hat{Q}_{i}=\mathrm{svec}^{-1}(\mathcal{E}(\Theta^{\dagger}_{T}\Psi_{T}\mathrm{svec}(\hat{P}_{i})+\Theta^{{\dagger}}_{T}\Xi_{T}))\,.$ (65)
@@
Since $\hat{P}_{0}\in\mathcal{P}(\hat{P}_{0})$, by Proposition 5, $\left\|\Delta Q_{1}\right\|_{2}<d^{*}$ almost surely. By definition, $\hat{P}_{1}\in\mathcal{P}(\hat{P}_{0})$. Thus, by induction, we have $\left\|\Delta Q\right\|_{\infty}<d^{*}$ almost surely. By Theorem 4, the result follows. ∎
@@ 5. Empirical Results
  • Nominal Control [50, 18]: a plugin PI (VI) model-based procedure that first estimates the system and cost matrices $(A,B,Q,R)$ and then solves the stochastic LQR problem using Hewer’s PI [23] (Bertsekas’s VI see [7]).
@@
This distinction matters for value iteration because the controller is derived from $\hat{P}$ (computed via the Lyapunov equation (75)) rather than from $\hat{P}_{{I_{\text{max}}}}$.
@@
We initialize each procedure with $x_{\text{init}}=0_{3}$ and $\hat{P}_{0}=\beta I_{3}$, with $\beta\sim\mathcal{U}\left(0,1\right)$ to ensure that $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$. Since the nominal PI, LSPI and O-LSPI are PI-based algorithms, it is necessary to select an initial control gain, $\hat{K}_{0}$. For consistency and fairness across comparisons, we choose $\hat{K}_{0}$ as the corresponding induced gains $\hat{K}_{0}=(R+B^{\top}\hat{P}_{0}B)^{-1}B^{\top}\hat{P}_{0}A$, as in [52]. For the policy gradient specifically, we take $\hat{K}_{0}$ stabilizing, $\hat{K}_{0}=\alpha_{0}I_{3}$, with $\alpha_{0}\sim\mathcal{U}\left(0.1,1\right)$ stabilizing to enhance its convergence.
@@ 5.2. Adaptivity under Non-Quadratic Costs
$\tilde{c}(x_{t},u_{t})\triangleq x_{t}^{\top}Sx_{t}+(|u_{t}|^{\frac{{\kappa}}{2}})^{\top}R|u_{t}|^{\frac{{\kappa}}{2}}$ (77)
where $|\cdot|^{\frac{{\kappa}}{2}}$ is applied elementwise and ${\kappa}\in[1,3]$. Since this cost function is non-quadratic in $u$ for values other than ${\kappa}=2$, the optimal cost is not expected to adhere to the standard form $\mathrm{Tr}\left(C^{\top}P^{*}C\right)$. Consequently, direct comparisons of relative costs are not meaningful in this setting. Instead, we evaluate the cost obtained using a control input $u_{t}=-\hat{K}x_{t}$, where $\hat{K}$ is computed for each method using a sample size of $T=10^{5}$. The choice of $T$ is guided by prior results (Figure 1), which show that this sample size enables most methods to reach $100\%$ stability. To assess adaptability to cost misspecification independently of the non-stabilizing data-collection policy, we set $K_{c}=\alpha I_{3}$ with $\alpha\sim\mathcal{U}(0.1,0.2)$.
Figure 2. Convergence and stability of the nominal control (VI and PI), LSPI, O-LSPI, R-LSVI and policy gradient algorithms on the data center cooling benchmark problem with non-quadratic cost. (Left) Costs as the exponent $\kappa$} varies. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
In Figure 2, we observe that the ADP algorithms consistently outperform nominal model-based control in both adaptability and stability. In the left panel, for ${\kappa}=2$, the nominal control slightly outperforms the other methods, which is consistent with the exact model specification and the findings in Figure 1. However, beyond this point, R-LSVI systematically outperforms all other algorithms, exhibiting superior adaptability. Notably, the right panel shows that for ${\kappa}>2$, nominal model-based control becomes unstable, with stability dropping sharply below $20\%$. In contrast, non-model-based methods maintain robust performance, with R-LSVI achieving the highest stability, remaining above $80\%$ across all values of $\kappa$}.